OpenStudy (anonymous):

solve x^4=1 for x should be 4 solutions. lol

OpenStudy (anonymous):

x=1, x=-1

OpenStudy (anonymous):

i guess 2 only

OpenStudy (anonymous):

(x+1)(x-1)(x+1)(x-1)?

OpenStudy (akshay_budhkar):

no x=1 and -1 are the only solution.

OpenStudy (akshay_budhkar):

yes the thing you did is right

OpenStudy (akshay_budhkar):

that is 0 so only x=1 or x=-1 done

OpenStudy (anonymous):

so am i right???

OpenStudy (akshay_budhkar):

lol yea!! don't doubt that nick! you are

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{w}}\color{orange}{\text{h}}\color{#9c9a2e}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{i}}\color{purple}{\text{s}}\color{purple}{\text{ }}\color{red}{\text{r}}\color{orange}{\text{i}}\color{#9c9a2e}{\text{g}}\color{green}{\text{h}}\color{blue}{\text{t}}\color{purple}{\text{?}}\color{purple}{\text{?}}\color{red}{\text{}}\end{array}\]

OpenStudy (anonymous):

-1and -1 is 1answer...so only 2solution exist...

jimthompson5910 (jim_thompson5910):

x^4=1 x^4-1=0 (x^2)^2-1^2=0 (x^2-1)(x^2+1)=0 (x-1)(x+1)(x^2+1)=0 x-1=0, x+1=0, or x^2+1=0 x=1, x=-1, or x^2=-1 x=1, x=-1, x=sqrt(-1) or x=-sqrt(-1) x=1, x=-1, x=i or x=i

OpenStudy (anonymous):

my teacher is VERY smart. I think there are four lol.

jimthompson5910 (jim_thompson5910):

oops meant to say x=1, x=-1, x=i or x=-i

jimthompson5910 (jim_thompson5910):

yes there are 4 and those are the solutions

jimthompson5910 (jim_thompson5910):

rule of thumb: degree of polynomial = number of solutions

jimthompson5910 (jim_thompson5910):

of course, I'm talking about complex solutions

OpenStudy (anonymous):

there is a reason I'm a fan of you jim.

OpenStudy (anonymous):

wait could I say (x²-1)(x²+1) = 0 and then x²=±1 and x²=±i?

jimthompson5910 (jim_thompson5910):

I think you mean x=±1 or x=±i, which is the shorthand for saying x=1, x=-1, x=i or x=-i

OpenStudy (anonymous):

oo, yeah, woops. lol

jimthompson5910 (jim_thompson5910):

Remember, the plus/minus only comes from taking the square root of both sides.

OpenStudy (anonymous):

even if one side is -?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

well x²= -1 does that mean x=±1?

OpenStudy (akshay_budhkar):

no it means +-i

jimthompson5910 (jim_thompson5910):

no, taking the square root of both sides means that x^2=-1 x=sqrt(-1) or x=-sqrt(-1) x=i or x=-i since i = sqrt(-1)

OpenStudy (anonymous):

oh, yeah that was my mistake but you still answered the intent of my question lol

jimthompson5910 (jim_thompson5910):

that's good that you're asking though

OpenStudy (akshay_budhkar):

So you wanted complex roots to? nice job jim